OpenStudy (anonymous):

Solve.Find three consecutive odd numbers whose sum is 45.

OpenStudy (anonymous):

\[x+1+x+3+x+5=45\] \[3x+9=45\] \[3x=45-9\] \[3x=36\] \[x=36/3=12\]

OpenStudy (anonymous):

\[x+1=13\] \[x+3=15\] \[x+5=17\]

OpenStudy (lalaly):

x+(x+2)+(x+4)=45

OpenStudy (lalaly):

3x+6=45 3x=39 x=13 x+2=15 x+4=17